Can you help?

ON Company manufactures windows. Its manufacturing plant has the capacity to produce 12,000 windows each month. Current production and sales are 10,000 windows per month. The company normally charges $250 per window. Cost information for the current activity level is as follows:
Variable Costs (per unit)
Direct Material $60
Direct Manufacturing labour $75
Variable Costs (material handling, quality control, etc) $15
Fixed manufacturing costs $250,000
Fixed marketing costs $400,000
ON has just received a special one-time-only order for 2,000 windows at $225 per window. Accepting the special order would not affect the company’s regular business or its fixed costs.
Prepare based on totals units and total revenue/costs  
Required:
1 Should ON accept this special order? What is the benefits to the company. Show your calculations.
2 Suppose plant capacity were only 10,000 windows instead of 12,000 windows each month. The special order must either be taken in full or be rejected completely. Should ON accept the special order? Why or Why not? Show your calculations.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
ON Company
Answer 1
To decide whether to accept this special order or not we have to calculate incremental income from the special order. The workings are below:
Incremental income statement
Amount $ Note
Sales volume            2,000.00 A
Sell Price               225.00 B
Less: Variable cost per unit
Direct Material                 60.00
Direct Manufacturing labor                 75.00
Other Variable Costs                 15.00
Total Variable cost per unit               150.00 C
Contribution margin per unit                 75.00 D=B-C
Total Contribution       150,000.00 E=A*D
Incremental income       150,000.00 See E
Accepting the special order will have incremental income of $ 150,000. So yes the order should be accepted.
What is the benefits to the company?
$ 150,000 is the benefits to the company.
Answer 2
To decide whether to accept this special order or not we will compare income in both the cases:
No special order Special order
Amount $ Amount $ Note
Sales volume         10,000.00        8,000.00 F
Sell Price               250.00           250.00 G
Less: Variable cost per unit
Direct Material                 60.00             60.00
Direct Manufacturing labor                 75.00             75.00
Other Variable Costs                 15.00             15.00
Total Variable cost per unit               150.00           150.00 See C
Contribution margin per unit               100.00           100.00 H=G-C
Total Contribution 1,000,000.00 800,000.00 I=F*H
Less: Fixed costs
Fixed manufacturing costs       250,000.00 250,000.00
Fixed marketing costs       400,000.00 400,000.00
Total Fixed costs       650,000.00 650,000.00 J
Add: Incremental income from special order 150,000.00 See E
Net operating income       350,000.00 300,000.00 K=I-J+E
Reduced by     50,000.00
Accepting the special order will reduce Net operating income by $ 50,000. So the order should not be accepted.